# Quantifying discontinuity

### v1

@@ 2. Background and preliminaries
$\mathrm{VR}_{<}(X;r)\longrightarrow\mathcal{P}_{<}(X;r)$ (1)
@@
$\mathcal{P}(S^{n};r)\simeq\begin{cases}S^{n},&r<r_{n}\\ S^{n}*\frac{\mathrm{SO}(n+1)}{A_{n+2}},&r=r_{n},\end{cases}$ (2)
@@
The result of Dubins and Schwarz [9] mentioned above was generalized in the paper [3] by the present authors and collaborators:
Theorem 2.1 (Theorems 1.3 and 7.6 of [3]).
Any odd function $f\colon S^{k}\rightarrow S^{n}$ with $k\geq n$ has modulus of discontinuity $\delta(f)\geq c_{n,k}$, and this bound is tight.
We remark that Lim, Mémoli, and Smith [18] used Dubins and Schwarz’ work on the modulus of discontinuity [9] in order to lower bound the Gromov–Hausdorff distance between spheres of different dimensions. The existence of odd maps $S^{k}\to\mathrm{VR}(S^{n};r)$ was studied in Adams, Bush, and Frick [4], and related to coverings and packings in projective spaces. These papers were the motivation for defining the constants $c_{n,k}$ in [3], which furthered the study of Gromov–Hausdorff distances between spheres.
@@
$\inf\{r\geq 0\colon~\exists~\text{continuous~}X\to_{G}\mathrm{VR}(Y;r)\}=\inf\{r\geq 0\colon~\exists~\text{continuous~}X\to_{G}\mathcal{P}(Y;r)\}.$
Indeed, the infimum on the left hand side is not smaller due to the $G$-map $\mathrm{VR}(X;r)\to\mathcal{P}(X;r)$ induced by the identity. As for the other inequality, if there is a $G$-map $f\colon~X\to\mathcal{P}(Y;r)$, it can be lifted (up to $G$-equivariant homotopy) to a $G$-map $\widetilde{f}\colon~X\to\mathrm{VR}_{<}(Y;r+\varepsilon)\subseteq\mathrm{VR}(Y;r+\varepsilon)$, for any $\varepsilon>0$, as depicted in the following diagram of $G$-maps (that commutes up to $G$-homotopy):
@@
For $G$-spaces AA and $B$, we will denote by $A,B]_{G}$ the set of $G$-maps $A\to B$ up to $G$-equivariant homotopy equivalence. Then, the composition of the two horizontal arrows in the diagram defines an element $f]\in[X,\mathcal{P}_{<}(Y;r+\varepsilon)]_{G}$. By [24, Proposition II.2.6], the vertical arrow in the diagram, which is a weak homotopy equivalence and a $G$-map, induces a surjection
@@
which shows that there is an element $\widetilde{f}]\in[X,\mathrm{VR}_{<}(Y;r+\varepsilon)]_{G}$ mapping to $[f]$.
@@
It is monotone in the sense that if there exists a $\mathbb{Z}/2$-map $X\to Y$, then $\mathrm{coind}_{\mathbb{Z}/2}(X)\leq\mathrm{coind}_{\mathbb{Z}/2}(Y)$. In general, if $X$ is (k1)(k-1)-connected, then $\mathrm{coind}_{\mathbb{Z}/2}(X)\geq k$. Moreover, Borsuk–Ulam theorem yields that $\mathrm{coind}_{\mathbb{Z}/2}(S^{k})=k$, where $\mathbb{Z}/2$ is assumed to act antipodally on $S^{k}$. We refer the reader to Matoušek’s book [20, Section 5] for more details. In the language of coindex, we can write the constants $c_{n,k}$ as $\inf\{r\colon~\mathrm{coind}_{\mathbb{Z}/2}(\mathrm{VR}(S^{n};r)\geq k)\}$.
@@ 3.1. Scale-invariant modulus of discontinuity $\alpha$
$\mathrm{Conf}_{2}(X)=\{(x,y)\in X\times X\ :\ x\neq y\}$
@@
$\Phi_{f}\colon\mathrm{Conf}_{2}(X)\longrightarrow S^{d-1},~(x,y)\longmapsto\frac{f(x)-f(y)}{\|f(x)-f(y)\|}.$ (3)
@@
$\mathrm{Conf}_{2}(X)\longrightarrow\mathcal{P}(S^{d-1};\delta(\Phi_{f})+\varepsilon).$ (4)
@@
By the assumption on coindex, there exists a $\mathbb{Z}/2$-map $g\colon~S^{k}\to\mathrm{Conf}_{2}(X)$. Note the composition $\Phi_{f}\circ g\colon S^{k}\to S^{d-1}$ is a $\mathbb{Z}/2$-map. Therefore by Theorem 2.1, $\delta(\Phi_{f}\circ g)\geq c_{d-1,k}$. Finally, we obtain the claim since $\alpha(f)=\delta(\Phi_{f})\geq\delta(\Phi_{f}\circ g)$ by [3, Lemma 7.10(ii)]. ∎
@@ 3.2. Suitability of $\alpha$ for detecting discontinuity
$\left\langle\frac{f(x)-f(y)}{\|f(x)-f(y)\|},\frac{f(x_{n})-f(y)}{\|f(x_{n})-f(y)\|}\right\rangle<1-\varepsilon$ (5)
@@ 3.3. Van Kampen–Flores theorem
$P_{f}\colon\mathrm{Conf}_{2}^{\Delta}(K)\longrightarrow S^{2d-1},~(x,y)\longmapsto\frac{f(x)-f(y)}{\|f(x)-f(y)\|},$ (6)
@@
$\mathrm{Conf}_{2}^{\Delta}(\operatorname{\mathrm{sk}}_{d}(\Delta_{2d+2}))\longrightarrow\mathcal{P}(S^{2d-1};\alpha^{(2)}(f)+\varepsilon).$ (7)
@@
Let us introduce the 2-fold deleted join (see [20, Section 5.5] for more detail) as a simplicial complex
@@
However, this contradicts [20, Section 5.6, pg. 117] which says that $(\operatorname{\mathrm{sk}}_{d}(\Delta_{2d+2}))^{*}_{\Delta}$ is a $(2d+1)$-sphere. ∎
@@
exists. Namely, the domain is a $(2d)$-dimensional cellular complex with a free $\mathbb{Z}/2$-action and by (2) the codomain is (2d1)(2d-1)-connected. Hence, all obstructions to defining such a $\mathbb{Z}/2$-map vanish (see also [20, Lemma 6.2.2]).

### v2

@@ Abstract.
2020 Mathematics Subject Classification. 51F30, 57N35, 52A35.
HA was supported by the Simons Foundation Travel Support for Mathematicians. FF was supported by NSF CAREER Grant DMS 2042428. NS was funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy–The Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID 390685689, BMS Stipend).
@@ 2. Background and preliminaries
(1) $\mathrm{VR}_{<}(X;r)\longrightarrow\mathcal{P}_{<}(X;r)$
@@
(2) $\mathcal{P}(S^{n};r)\simeq\begin{cases}S^{n},&r<r_{n}\\ S^{n}*\frac{\mathrm{SO}(n+1)}{A_{n+2}},&r=r_{n},\end{cases}$
@@
The result of Dubins and Schwarz [9] mentioned above was generalized by the present authors and collaborators in [3, Theorems 1.3 and 7.6], where it is proved that for $k\geq n$, any odd function $f\colon S^{k}\rightarrow S^{n}$ has modulus of discontinuity
(3) $\delta(f)\geq c_{n,k}$
and that this bound is tight. We remark that Lim, Mémoli, and Smith [19] used Dubins and Schwarz’ work on the modulus of discontinuity [9] in order to lower bound the Gromov–Hausdorff distance between spheres of different dimensions. The existence of odd maps $S^{k}\to\mathrm{VR}(S^{n};r)$ was studied in Adams, Bush, and Frick [4], and related to coverings and packings in projective spaces. These papers were the motivation for defining the constants $c_{n,k}$ in [3], which furthered the study of Gromov–Hausdorff distances between spheres.
@@
(4) $\inf\{r\geq 0\colon~\exists~\text{continuous~}X\to_{G}\mathrm{VR}(Y;r)\}=\inf\{r\geq 0\colon~\exists~\text{continuous~}X\to_{G}\mathcal{P}(Y;r)\}.$
We point out that the constant on the left hand side was described to us by Facundo Mémoli in a Polymath meeting in 2022 (see also [18, Definition 24]). Indeed, the infimum on the left hand side is not smaller due to the $G$-map $\mathrm{VR}(X;r)\to\mathcal{P}(X;r)$ induced by the identity. As for the other inequality, if there is a $G$-map $f\colon~X\to\mathcal{P}(Y;r)$, it can be lifted (up to $G$-equivariant homotopy) to a $G$-map $\widetilde{f}\colon~X\to\mathrm{VR}_{<}(Y;r+\varepsilon)\subseteq\mathrm{VR}(Y;r+\varepsilon)$, for any $\varepsilon>0$, as depicted in the following diagram of $G$-maps (that commutes up to $G$-homotopy):
@@
For $G$-spaces AA and $B$, we will denote by $A,B]_{G}$ the set of $G$-maps $A\to B$ up to $G$-equivariant homotopy equivalence. Then, the composition of the two horizontal arrows in the diagram defines an element $f]\in[X,\mathcal{P}_{<}(Y;r+\varepsilon)]_{G}$. By [25, Proposition II.2.6], the vertical arrow in the diagram, which is a weak homotopy equivalence and a $G$-map, induces a surjection
@@
which shows that there is an element $\widetilde{f}]\in[X,\mathrm{VR}_{<}(Y;r+\varepsilon)]_{G}$ mapping to $[f]$. The equality (4) was obtained in a different way by explicitly constructing maps between Vietoris-Rips complexes and metric thickenings (see [18, Proposition 6.2]). It is also possible to obtain a generalization of (3) using the quantities in (4) to the case when $G$ is a finite group and $f\colon X\to Y$ is any $G$-function between $G$-metric spaces. We do this in special cases of interest (see proofs of Theorems 3.3, 3.11, and 4.3); for a general case, see [18, Proposition 6.5].
Our paper and the recent paper of Lim and Mémoli [18] are follow-up projects to the polymath-style paper [3]; as such, the two papers use related methods towards different goals. The main goal of [18] is to define equivariant distances, obtain bounds, and establish equivariant rigidity results.
@@
It is monotone in the sense that if there exists a $\mathbb{Z}/2$-map $X\to Y$, then $\mathrm{coind}_{\mathbb{Z}/2}(X)\leq\mathrm{coind}_{\mathbb{Z}/2}(Y)$. In general, if $X$ is (k1)(k-1)-connected, then $\mathrm{coind}_{\mathbb{Z}/2}(X)\geq k$. Moreover, Borsuk–Ulam theorem yields that $\mathrm{coind}_{\mathbb{Z}/2}(S^{k})=k$, where $\mathbb{Z}/2$ is assumed to act antipodally on $S^{k}$. We refer the reader to Matoušek’s book [21, Section 5] for more details. In the language of coindex, we can write the constants $c_{n,k}$ as $\inf\{r\colon~\mathrm{coind}_{\mathbb{Z}/2}(\mathrm{VR}(S^{n};r)\geq k)\}$.
@@ 3.1. Scale-invariant modulus of discontinuity $\alpha$
$\mathrm{Conf}_{2}(X)=\{(x,y)\in X\times X\colon~x\neq y\}$
@@
(5) $\Phi_{f}\colon\mathrm{Conf}_{2}(X)\longrightarrow S^{d-1},~(x,y)\longmapsto\frac{f(x)-f(y)}{\|f(x)-f(y)\|}.$
@@
(6) $\mathrm{Conf}_{2}(X)\longrightarrow\mathcal{P}(S^{d-1};\delta(\Phi_{f})+\varepsilon).$
@@
By the assumption on coindex, there exists a $\mathbb{Z}/2$-map $g\colon~S^{k}\to\mathrm{Conf}_{2}(X)$. Note the composition $\Phi_{f}\circ g\colon S^{k}\to S^{d-1}$ is a $\mathbb{Z}/2$-map. Therefore by (3), $\delta(\Phi_{f}\circ g)\geq c_{d-1,k}$. Finally, we obtain the claim since $\alpha(f)=\delta(\Phi_{f})\geq\delta(\Phi_{f}\circ g)$ by [3, Lemma 7.10(ii)]. ∎
@@ 3.2. Suitability of $\alpha$ for detecting discontinuity
(7) $\left\langle\frac{f(x)-f(y)}{\|f(x)-f(y)\|},\frac{f(x_{n})-f(y)}{\|f(x_{n})-f(y)\|}\right\rangle<1-\varepsilon$
@@ 3.3. Van Kampen–Flores theorem
(8) $P_{f}\colon\mathrm{Conf}_{2}^{\Delta}(K)\longrightarrow S^{2d-1},~(x,y)\longmapsto\frac{f(x)-f(y)}{\|f(x)-f(y)\|},$
@@
(9) $\mathrm{Conf}_{2}^{\Delta}(\operatorname{\mathrm{sk}}_{d}(\Delta_{2d+2}))\longrightarrow\mathcal{P}(S^{2d-1};\alpha^{(2)}(f)+\varepsilon).$
@@
Let us introduce the 2-fold deleted join (see [21, Section 5.5] for more detail) as a simplicial complex
@@
However, this contradicts [21, Section 5.6, pg. 117] which says that $(\operatorname{\mathrm{sk}}_{d}(\Delta_{2d+2}))^{*}_{\Delta}$ is a $(2d+1)$-sphere. ∎
@@
exists. Namely, the domain is a $(2d)$-dimensional cellular complex with a free $\mathbb{Z}/2$-action and by (2) the codomain is (2d1)(2d-1)-connected. Hence, all obstructions to defining such a $\mathbb{Z}/2$-map vanish (see also [21, Lemma 6.2.2]).